# OccLE: Label-Efficient 3D Semantic Occupancy Prediction

### v1

@@ Abstract
3D semantic occupancy prediction offers an intuitive and efficient scene understanding and has attracted significant interest in autonomous driving perception. Existing approaches either rely on full supervision, which demands costly voxel-level annotations, or on self-supervision, which provides limited guidance and yields suboptimal performance. To address these challenges, we propose OccLE, a Label-Efficient 3D Semantic Occupancy Prediction that takes images and LiDAR as inputs and maintains high performance with limited voxel annotations. Our intuition is to decouple the semantic and geometric learning tasks and then fuse the learned feature grids from both tasks for the final semantic occupancy prediction. Therefore, the semantic branch distills 2D foundation model to provide aligned pseudo labels for 2D and 3D semantic learning. The geometric branch integrates image and LiDAR inputs in cross-plane synergy based on their inherency, employing semi-supervision to enhance geometry learning. We fuse semantic-geometric feature grids through Dual Mamba and incorporate a scatter-accumulated projection to supervise unannotated prediction with aligned pseudo labels. Experiments show that OccLE achieves competitive performance with only 10% of voxel annotations, reaching a mIoU of 16.59% on the SemanticKITTI validation set.
@@
Keywords Autonomous Driving  $\cdot$ 3D Semantic Occupancy Prediction  $\cdot$ Label-Efficient Learning
1 Introduction
3D perception task is foundational cornerstone for autonomous driving systems. Among various perception methods, 3D semantic occupancy prediction [1, 2, 3] has garnered significant attention for providing intuitive and efficient scene understanding for downstream tasks. This task aims to estimate the occupancy status and semantic label of each voxel in a 3D grid, given input from 2D images, 3D LiDAR, or a combination of both.
Previous studies [4, 5] proposed numerous supervised learning paradigms with diverse modal inputs. A notable challenge is that their high performance relies heavily on extensive voxel annotations, which are both costly and labor-intensive. The cubic complexity of the voxel grid leads to a substantial workload during manual annotation. Even with pre-annotated 3D labels generated using auto-labeling assistants, one hour of human labor just only annotates 10 frames [6, 7]. This issue limits the scalability of supervised methods for real-world deployment, especially in terms of robustness.
@@ 1 Introduction
Recent advancements in 3D semantic occupancy prediction have shifted towards a label-free paradigm. These self-supervised methods leverage existing vision foundation models to generate various pseudo labels for auxiliary supervision, such as image semantic segmentation [8], depth information [9], and LiDAR segmentation [10]. To bridge the 3D representation with 2D pseudo label, specific 3D representation formats have been proposed to facilitate volume rendering, including signed distance functions[7] and 3D Gaussians [11]. The core strategy of these self-supervised methods is to supplement pseudo semantic supervision. They project LiDAR segmentation annotations from the same dataset into 2D pseudo labels [12, 6]. However, this pseudo labels generation still is expensive and provides only sparse supervision. They also employ a 2D vision foundation model to generate 2D pseudo labels [7], which suffers from class misalignment. For the overall pipeline, they supervise only the final prediction using 2D pseudo labels and do not apply separate supervision to geometric and semantic feature learning. Thus, this non-decoupled design leads to insufficient geometric learning, which reduces volume-rendering quality and weakens semantic supervision.
@@
To address these challenges, we propose OccLE, a label-efficient 3D semantic occupancy prediction method, as illustrated in Fig. 1, which uses limited voxel annotations while maintaining high performance. We design the label-efficient learning paradigm: taking image and LiDAR scan inputs, fully decoupling semantic and geometric feature learning, fusing their feature grids for 3D semantic occupancy prediction, and supervising with limited voxel annotations and aligned pseudo labels. Specifically, OccLE consists of the semantic branch, the geometric branch, and semantic-geometric feature grid fusion. The semantic branch distills dataset-specific and open-vocabulary 2D foundation models to produce aligned pseudo labels for supervising 2D and 3D semantic learning, independently of geometry. The geometric branch integrates image and LiDAR and inputs in cross-plane synergy based on their inherent characteristics, employing semi-supervision to enhance geometry learning. The semantic-geometric feature grid fusion employs Dual Mamba to fuse the two feature grids for lightweight long-range relationship modeling, and introduces a scatter-accumulated projection to guide unannotated predictions with aligned pseudo labels, without relying on any specific feature grid format. In summary, the main contributions of this work are summarized as follows: In summary, the main contributions of this work are summarized as follows:
@@
• We introduce a label-efficient 3D semantic occupancy prediction task and a learning paradigm that decouples semantic and geometric learning, followed by their synergistic integration. The model is supervised using limited voxel annotations and aligned pseudo labels to achieve high performance.
@@
• We present three techniques: distilling 2D foundation model for 2D and 3D semantic learning; semi-supervised geometry learning using cross-plane image and LiDAR feature synergy; and semantic-geometric feature grids fusion with auxiliary supervision by scatter-accumulated projection.
@@
• Experimental results show that OccLE achieves competitive performance compared to fully supervised methods, attaining 16.59 % mIoU on the SemanticKITTI validation set using only 10 % of the voxel annotations.
@@ 2 Related Works
For camera-based 3D semantic occupancy prediction, a key challenge lies in lifting 2D to 3D representations. Previous works [13, 14, 15] utilize calibration and depth information to establish explicit backwards projections. ViewFormer [16] proposes a learning-first attention mechanism to aggregate multi-view features. Some studies [17, 18] represent scenes using sparse 3D semantic Gaussians. TPVFormer [19] introduces three perpendicular views (TPV) to enhance the interaction of image features.
Since dimensional transformation based on depth estimation is ill-posed for camera-based methods, multi-modal method incorporates LiDAR and radar inputs to obtain accurate distance measurements. Its essential research is modal fusion. OccFusion [5] directly concatenates feature channels. MetaOcc [20] learn global and local alignment to enhance fused representations. Some studies [21, 22] advance fusion by employing point-to-point queries from 3D to 2D. OccGen [23] employs a diffusion model to refine the fused features.
Building on prior work, OccLE takes camera images and a LiDAR scans as inputs, applies backward projection to derive 3D feature grids from 2D image feature maps, and proposes a lightweight yet effective fusion strategy in the geometry learning termed cross-plane image and LiDAR feature synergy.
@@ 2.2 Label-efficient learning
In label-efficient learning for 3D semantic occupancy prediction, self-supervision methods render 3D feature into 2D representations, and supervise with labels from other tasks. To facilitate rendering, SelfOcc [7] introduces a signed distance function-based representation, while GaussTR [11] proposes a 3D Gaussian-based representation. For auxiliary supervision, OccFlowNet [12] utilizes projected semantic point clouds, whereas RenderOcc [6] relies on depth estimation and image segmentation.
@@ 3.1 Overview
We aim to take multi-frame camera images ${\left\{{{{\bf{I}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{h\times w}}}$, multi-frame point clouds ${\left\{{{{\bf{P}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{n\times 4}}}$, camera calibration ${\left\{{{\bf{K}}{\rm{,}}{\bf{T}}}\right\}}$ as inputs to predict the 3D semantic occupancy ${{{\bf{Y}}_{t}}\in{\left\{{{c_{0}},{c_{1}},\cdots,{c_{M-1}}}\right\}^{H\times W\times Z}}}$ at timestep ${t}$. To enable label-efficient training, we supervise the model using only limited voxel annotations (e.g., 10% ${\left\{{{{{\bf{\bar{Y}}}}_{t}}}\right\}}$ of all samples). Here, ${{\bf{K}}{\rm{,}}{\bf{T}}}$ represent the intrinsic and extrinsic camera matrices, respectively; ${\left\{{{{{\bf{\bar{Y}}}}_{t}}}\right\}}$ is the ground truth of 3D semantic occupancy prediction. ${N}$ denotes frame number; ${M}$ denotes class number, ${h}$ and ${w}$ denote the height and width of camera image; ${H}$, ${W}$, and ${Z}$ represent the length, width, height of voxel grid. Our method is formulated as ${{{\bf{Y}}_{t}}=\Theta\left({\left\{{{{\bf{I}}_{t-i}}}\right\}_{i=0}^{N-1},\left\{{{{\bf{P}}_{t-i}}}\right\}_{i=0}^{N-1},{\bf{K}}{\rm{,}}{\bf{T}}}\right)}$, where ${\Theta}$ is our proposed OccLE. For clarity, we omit the timestep subscript ${t}$ for all variables in the following description.
The overview of OccLE is illustrated in Fig. 2. For simplicity, only a single input frame is shown. To enable label-efficient learning, our pipeline comprises three components: a semantic branch, a geometric branch, and a semantic-geometric feature grid fusion module. We decouple the semantic and geometric learning with distinct supervision strategies, and fuse their feature grids to predict 3D semantic occupancy. In the semantic branch, we distill 2D foundation models to predict aligned pseudo labels and supervise 2D and 3D semantic feature learning. In the geometric branch, we synergize image and LiDAR features based on their inherency and employ semi-supervision to strengthen geometry learning. In the semantic-geometric feature grid fusion module, we employ Dual Mamba to fuse feature grids and employ scatter-accumulated projection to supervise unanotated predictions with aligned pseudo labels. The semantic branch, the geometric branch, and the semantic-geometric feature grid fusion will be discussed in Sec. 3.2, Sec. 3.3, and Sec. 3.4, respectively.
@@ 3.2 Distill 2D Foundation Models for 2D and 3D Semantic Learning
3D semantic occupancy prediction is implemented through the entanglement of geometric and semantic features [4]. The geometric feature indicates whether a voxel is occupied, while the semantic feature specifies the class occupying the voxel. 3D semantic occupancy prediction maps the feature grid to voxel semantic labels. Achieving high accuracy traditionally requires extensive voxel annotations to guide the mapping, but obtaining such annotations is costly. To mitigate this requirement, we notice that the image semantic segmentation maps the feature map to pixel semantic label and benefits from large annotated datasets. Therefore, we aim to distill 2D foundation model for 2D and 3D semantic learning, where the 2D foundation model predicts aligned pseudo labels to supervise the semantic learning of image and LiDAR scan inputs.
@@
We employ ${{\bf{\bar{s}}}}$ to fully supervise semantic learning of images and LiDAR scans, enabling lightweight deployment via distillation. Specifically, we employ a UNet [29] ${{{\cal G}_{s2d}}}$ to extract features map ${{\bf{F}}_{I}^{s}}$ from the image ${{\bf{I}}}$. We adopt the voxelization in VoxelNet [30] and utilize a sparse UNet3D [31] ${{{\cal G}_{s3d}}}$ to extract features grid ${{\bf{F}}_{P}^{s}}$ from the LiDAR scan ${{\bf{P}}}$. The mathematical expression is given by:
[…]
@@
where ${{\cal F}}$ denotes the sampling function, ${{\bf{P}}}$ denote the voxel grid coordinates, ${{\bf{p}}}$ represent the corresponding projected coordinates on the image plane.
@@ 3.3 Semi-Supervised Geometry Learning
Prior works [4, 32] depend on absolute depth in learning geometry for 3D semantic occupancy prediction. Because depth estimation is ill-posed, reducing voxel annotations exacerbates uncertainty in this sequential prediction. Since depth is sourced from the LiDAR scan projected on the image plane, we directly use image and LiDAR scan inputs, propose a lightweight and modality-complementary geometry learning module, termed cross-plane image and LiDAR feature synergy, and employ a semi-supervision to learn an accurate geometric feature grid with limited voxel annotations.
@@ 3.3.1 Cross-Plane Image and LiDAR Feature Synergy
First, we examine the characteristics of the image and LiDAR scan inputs. As shown by the red boxes in Fig. 3a and Fig. 3b, the frontal-view LiDAR scan is sparse and omits information in the upper region, whereas the image offers continuous context. In the bird’s-eye view (BEV), the image cannot resolve distances along each ray, while the LiDAR scan provides accurate range measurements. In summary, the image is reliable on the ${yz}$-plane and the LiDAR scan is reliable along the ${x}$-axis in world coordinates. Motivated by their inherency and inspired by [33], we project each their feature grid into its optimal plane in the TPV and fuse them via multiplicative integration.
@@
The cross-plane image and LiDAR feature synergy is illustrated in Fig. 3c. We voxelize multi-frame LiDAR scans using the VoxelNet [30] and process them with a sparse 3D encoder to obtain LiDAR feature grids. In parallel, we extract feature maps from multi-frame images using a 2D encoder and lift into the feature grid defined in Equ. 2. We fuse the multi-frame feature grids into a single grid via a Conv1D layer. For the LiDAR feature grid, we stack along the ${y}$ and ${z}$ axes and apply a Conv2D layer to produce feature maps ${{\bf{F}}_{I,xz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{Z_{1}}}}}$ and ${{\bf{F}}_{I,xy}^{g}\in{{\mathbb{R}}^{{W_{1}}\times{Z_{1}}}}}$. For the image feature grid, we stack along the ${x}$-axis and apply a Conv2D layer to generate feature map ${{\bf{F}}_{P,yz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$. To enhance ${{\bf{F}}_{P,xy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$, we apply self-attention (SA) along the ${y}$ and ${z}$ axes, yielding feature maps ${{\bf{F}}_{P,yzy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$, ${{\bf{F}}_{P,yzz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}}$ as follows:
@@
We add ${{\bf{F}}_{{f_{1}}}^{g}}$ and ${{\bf{F}}_{{f_{2}}}^{g}}$ to generate the geometric feature grid ${{\bf{F}}_{{f}}^{g}}$ and predict geometry ${{\bf{G}}\in{{\mathbb{R}}^{H\times W\times Z}}}$ via a 3D header.

### v2

@@ Abstract
3D semantic occupancy prediction offers an intuitive and efficient scene understanding and has attracted significant interest in autonomous driving perception. Existing approaches either rely on full supervision, which demands costly voxel-level annotations, or on self-supervision, which provides limited guidance and yields suboptimal performance. To address these challenges, we propose OccLE, a Label-Efficient 3D Semantic Occupancy Prediction that takes images and LiDAR as inputs and maintains high performance with limited voxel annotations. Our intuition is to decouple the semantic and geometric learning tasks and then fuse the learned feature grids from both tasks for the final semantic occupancy prediction. Therefore, the semantic branch distills 2D foundation model to provide aligned pseudo labels for 2D and 3D semantic learning. The geometric branch integrates image and LiDAR inputs in cross-plane synergy based on their inherency, employing semi-supervision to enhance geometry learning. We fuse semantic-geometric feature grids through Dual Mamba and incorporate a scatter-accumulated projection to supervise unannotated prediction with aligned pseudo labels. Experiments show that OccLE achieves competitive performance with only 10% of voxel annotations on the SemanticKITTI and Occ3D-nuScenes datasets.
Keywords— Autonomous Driving 3D Semantic Occupancy Prediction Label-Efficient Learning
1 Introduction
3D perception task is foundational cornerstone for autonomous driving systems. Among various perception methods, 3D semantic occupancy prediction [36, 24] has garnered significant attention for providing intuitive and efficient scene understanding for downstream tasks. This task estimates the occupancy status and semantic label of each voxel in a 3D grid, given input from 2D images, LiDAR, or a combination of both.
Previous studies [22, 30] proposed numerous supervised learning paradigms with diverse modal inputs. A notable challenge is that their high performance relies heavily on extensive voxel annotations, which are both costly and labor-intensive. The cubic complexity of the voxel grid leads to a substantial workload during manual annotation. Even with pre-annotated 3D labels generated using auto-labeling assistants, one hour of human labor just only annotates 10 frames [31, 14]. This issue limits the scalability and robustness of supervised methods in real-world deployment.
@@ 1 Introduction
Recent advancements in 3D semantic occupancy prediction have shifted towards a label-free paradigm. These self-supervised methods leverage existing vision foundation models to generate various pseudo labels for auxiliary supervision, such as image semantic segmentation [32], depth information [48], and LiDAR segmentation [57]. To bridge the 3D representation with 2D pseudo label, they proposed specific 3D representation formats to facilitate volume rendering, including signed distance functions [14] and 3D Gaussians [18]. The core strategy of these self-supervised methods is to supplement pseudo semantic supervision. They project LiDAR segmentation annotations from the same dataset into 2D pseudo labels [3, 31]. However, this pseudo labels still is expensive and provides only sparse supervision. They also employ a 2D vision foundation model to generate 2D pseudo labels [14], which suffers from class misalignment. For the overall pipeline, they supervise only the final prediction using 2D pseudo labels and do not apply separate supervision to geometric and semantic feature learning. Thus, this non-decoupled design leads to insufficient geometric learning, which reduces volume-rendering quality and weakens semantic supervision.
@@
To address these challenges, we propose OccLE, a label-efficient 3D semantic occupancy prediction method, as illustrated in Fig. 1, which uses limited voxel annotations while maintaining high performance. We design the label-efficient learning paradigm: taking image and LiDAR inputs, fully decoupling semantic and geometric feature learning, fusing their feature grids for 3D semantic occupancy prediction, and supervising with limited voxel annotations and aligned pseudo labels. Specifically, OccLE consists of the semantic branch, the geometric branch, and semantic-geometric feature grid fusion. The semantic branch distills dataset-specific and open-vocabulary 2D foundation models to produce aligned pseudo labels for supervising 2D and 3D semantic learning, independently of geometry. The geometric branch integrates image and LiDAR and inputs in cross-plane synergy based on their inherent characteristics, employing semi-supervision to enhance geometry learning. The semantic-geometric feature grid fusion employs Dual Mamba to fuse the two feature grids for lightweight long-range relationship modeling, and introduces a scatter-accumulated projection to guide unannotated predictions with aligned pseudo labels, without relying on any specific feature grid format. In summary, the main contributions of this work are summarized as follows: In summary, the main contributions of this work are summarized as follows:
@@
• We introduce a label-efficient 3D semantic occupancy prediction task and a learning paradigm that decouples semantic and geometric learning, followed by their synergistic integration. The model is supervised with limited voxel annotations and aligned pseudo labels to achieve high performance.
@@
• We present three techniques: distilling 2D foundation model for 2D and 3D semantic learning; semi-supervised geometry learning using cross-plane image and LiDAR feature synergy; and semantic-geometric feature grids fusion with scatter-accumulated projection auxiliary supervision.
@@
• Experimental results show that OccLE achieves competitive performance compared to fully supervised methods, attaining 16.59 % and 27.53% mIoU on the SemanticKITTI validation set and Occ3D-nuScenes using only 10 % of the voxel annotations.
@@ 2 Related Works
For camera-based 3D semantic occupancy prediction, a key challenge lies in lifting 2D to 3D representations. Previous works [53, 42] utilize calibration and depth information to establish explicit backwards projections. ViewFormer [21] proposes a learning attention mechanism to aggregate multi-view features. Some studies [8, 16] represent scenes using 3D Gaussians. TPVFormer [13] introduces three perpendicular views (TPV) to enhance the interaction of image features.
Since dimensional transformation based on estimated depth is ill-posed for camera-based methods, multi-modal method incorporates LiDAR and radar inputs to obtain accurate distance measurements. Its essential research is modal fusion. OccFusion [30] directly concatenates feature channels. MetaOcc [46] learn global and local alignment to enhance fused representations. Some studies [50, 52] employ point-to-point queries from 3D to 2D. OccGen [38] employs a diffusion model to refine the fused features.
Building on prior work, OccLE takes camera images and LiDAR scans as inputs, applies backward projection to derive 3D feature grids from 2D image feature maps, and proposes a lightweight fusion strategy for geometry learning, termed cross-plane image and LiDAR feature synergy.
@@ 2.2 Label-efficient learning
In label-efficient learning for 3D semantic occupancy prediction, self-supervision methods render 3D feature into 2D format, and supervise with labels from other tasks. To facilitate rendering, SelfOcc [14] introduces a signed distance function-based representation, while GaussTR [18] proposes a 3D Gaussian-based representation. For auxiliary supervision, OccFlowNet [3] utilizes projected LiDAR semantic, whereas RenderOcc [31] relies on estimated depth and image segmentation.
@@ 3.1 Overview
We aim to take multi-frame camera images $\left\{{{{\bf{I}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{h\times w\times 3}}$}, multi-frame point clouds $\left\{{{{\bf{P}}_{t-i}}}\right\}_{i=0}^{N-1}\in{{\mathbb{R}}^{n\times 4}}$}, camera calibration $\left\{{{\bf{K}}{\rm{,}}{\bf{T}}}\right\}$} as inputs to predict the 3D semantic occupancy ${{\bf{Y}}_{t}}\in{\left\{{{c_{0}},{c_{1}},\cdots,{c_{M-1}}}\right\}^{H\times W\times Z}}$} at timestep t{t}. To enable label-efficient training, we supervise the model training with limited voxel annotations (e.g., 10% $\left\{{{{{\bf{\bar{Y}}}}_{t}}}\right\}$} of all samples). Here, ${\bf{K}}{\rm{,}}{\bf{T}}$} represent the intrinsic and extrinsic camera matrices, respectively; $\left\{{{{{\bf{\bar{Y}}}}_{t}}}\right\}$} is the ground truth of 3D semantic occupancy prediction. N{N} denotes frame number; M{M} denotes class number, h{h} and w{w} denote the height and width of camera image; H{H}, W{W}, and Z{Z} represent the length, width, height of voxel grid. Our method is formulated as ${{\bf{Y}}_{t}}=\Theta\left({\left\{{{{\bf{I}}_{t-i}}}\right\}_{i=0}^{N-1},\left\{{{{\bf{P}}_{t-i}}}\right\}_{i=0}^{N-1},{\bf{K}}{\rm{,}}{\bf{T}}}\right)$}, where $\Theta$} is our proposed OccLE. For clarity, we omit the timestep subscript t{t} for all variables in the following description.
The overview of OccLE is illustrated in Fig. 2. For simplicity, only a single input frame is shown. To enable label-efficient learning, our pipeline comprises three components: a semantic branch, a geometric branch, and a semantic-geometric feature grid fusion module. We decouple the semantic and geometric learning with distinct supervision strategies, and fuse their feature grids to predict 3D semantic occupancy. In the semantic branch, we distill 2D foundation models to predict aligned pseudo labels and supervise 2D and 3D semantic feature learning. In the geometric branch, we synergize image and LiDAR features based on their inherency and employ semi-supervision to strengthen geometry learning. In the semantic-geometric feature grid fusion module, we employ Dual Mamba to fuse feature grids and employ scatter-accumulated projection to supervise unanotated predictions with aligned pseudo labels.
@@ 3.2 Distill 2D Foundation Models for 2D and 3D Semantic Learning
3D semantic occupancy prediction is implemented through the entanglement of geometric and semantic features [22]. The geometric feature indicates whether a voxel is occupied, while the semantic feature specifies the class occupying the voxel. 3D semantic occupancy prediction maps the feature grid to voxel semantic labels. Achieving high accuracy traditionally requires extensive voxel annotations to guide the mapping, but obtaining such annotations is costly. To mitigate this requirement, we notice that the image semantic segmentation maps the feature map to pixel semantic label and benefits from large annotated datasets. Therefore, we aim to distill 2D foundation model for 2D and 3D semantic learning, where the 2D foundation model predicts aligned pseudo labels to supervise the semantic learning of image and LiDAR inputs.
@@
We employ $\bf{\bar{s}}$}} to fully supervise semantic learning of images and LiDAR, enabling lightweight deployment via distillation. Specifically, we employ a UNet [34] ${\cal G}_{s2d}$}} to extract features map ${\bf{F}}_{I}^{s}$} from the image $\bf{I}$}}. We adopt the voxelization in VoxelNet [56] and utilize a sparse UNet3D [44] ${\cal G}_{s3d}$}} to extract features grid ${\bf{F}}_{P}^{s}$} from the LiDAR $\bf{P}$}}. The mathematical expression is given by:
$\left\{\begin{aligned} \mathbf{F}_{I}^{s}&=\mathcal{G}_{s2}(\mathbf{I})\\ \mathbf{F}_{P}^{s}&=\mathcal{G}_{s3}(\mathcal{V}(\mathbf{P}))\end{aligned}\right$. (1)
@@
[…]
@@ 3.3 Semi-Supervised Geometry Learning
Prior works [22, 41] depend on absolute depth in learning geometry for 3D semantic occupancy prediction. Because depth estimation is ill-posed, reducing voxel annotations exacerbates uncertainty in this sequential prediction. Since depth is sourced from the LiDAR projected on the image plane, we directly use image and LiDAR inputs, propose a lightweight and modality-complementary geometry learning module, termed cross-plane image and LiDAR feature synergy, and employ a semi-supervision to learn an accurate geometric feature grid with limited voxel annotations.
@@ 3.3.1 Cross-Plane Image and LiDAR Feature Synergy
First, we examine the characteristics of the image and LiDAR inputs. As shown by the red boxes in Fig. 3a and Fig. 3b, the frontal-view LiDAR is sparse and omits information in the upper region, whereas the image offers continuous context. In the bird’s-eye view (BEV), the image cannot resolve distances along each ray, while the LiDAR provides accurate range measurements. In summary, the image is reliable on the yz{yz}-plane and the LiDAR is reliable along the x{x}-axis in world coordinates. Motivated by their inherency and inspired by [51], we propose using TPV to achieve a synergistic and lightweight multi-modal features fusion by projecting features onto their optimal TPV planes and applying multiplicative integration.
@@
The cross-plane image and LiDAR feature synergy is illustrated in Fig. 3c. We voxelize multi-frame LiDAR using the VoxelNet [56] and process them with a sparse 3D encoder to obtain LiDAR feature grids. In parallel, we extract feature maps from multi-frame images using a 2D encoder and lift into the feature grid defined in Equ. 2. We fuse the multi-frame feature grids into a single grid via a Conv1D layer. For the LiDAR feature grid, we stack along the y{y} and z{z} axes and apply a Conv2D layer to produce feature maps ${\bf{F}}_{I,xz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{Z_{1}}}}$} and ${\bf{F}}_{I,xy}^{g}\in{{\mathbb{R}}^{{W_{1}}\times{Z_{1}}}}$}. For the image feature grid, we stack along the x{x}-axis and apply a Conv2D layer to generate feature map ${\bf{F}}_{P,yz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}$}. To enhance ${\bf{F}}_{P,xy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}$}, we apply self-attention (SA) along the y{y} and z{z} axes, yielding feature maps ${\bf{F}}_{P,yzy}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}$}, ${\bf{F}}_{P,yzz}^{g}\in{{\mathbb{R}}^{{H_{1}}\times{W_{1}}}}$} as follows:
@@
We add ${\bf{F}}_{{f_{1}}}^{g}$} and ${\bf{F}}_{{f_{2}}}^{g}$} to obtain the geometric feature grid ${\bf{F}}_{{f}}^{g}$} and predict geometry ${\bf{G}}\in{{\mathbb{R}}^{H\times W\times Z}}$} via a 3D header.